OpenStudy (anonymous):

solve for P. p+40=2000-50P i got P=38.431?

OpenStudy (praxer):

p+50p=2000-40 =>51p=1960 =>p=1960/51 =>p=38.43

OpenStudy (praxer):

Cheers ! you got the correct answer.. :)

OpenStudy (anonymous):

Yes! Thanks!

OpenStudy (anonymous):

p +40 = 2000 - 50p 51p+40 = 2000 51p = 1960 1960/51 =38.4313725490196078...

OpenStudy (praxer):

Welcome.:)